use substitution to solve the system.
y=x-3
4x+y=1

Answer by Nate(3500)

For y=x-3 and 4x+y=1, you would substitute the answer for y in the second equation.
4x+y=1--original equation
4x+x-3=1--substitute y in for x
5x=4--simplify
x=4/5--divide whole equation by 5
Now, we have to find what y equals.
y=x-3--original equation
y=4/5-3--substitute x in for y
y=-11/5--simplify